OpenStudy (anonymous):

Kate and Jean paint a house together in 6 days. If they work alone, Kate takes 5 fewer days than Jean would take to do the same job. Which equation represents the time it takes Kate to paint the house by herself?

OpenStudy (anonymous):

the answer choices are: x^2−17x+30=0 x^2−5x−6=0 2^x−1=0 x^2−5x/6=0

OpenStudy (anonymous):

The first one i am pretty sure.